Telescope Status

Telescope Status Updates
26/04/2022 T2: colour camera bayer matrix corrected T11: requires new calibration frames T18: new QHY600M plus LRGB HSO V and I filters being...
Tue, 26 Apr, 2022 at 3:41 PM